When a flower shop raised the price of a floral arrangement from $20 to $28, the number of the arrangements sold decreased from 30 a week to 20. What is the price elasticity of demand for the flowers in this case?  


1
Expert's answer
2021-01-31T19:30:53-0500

Price elasticity of demand (PED) is the percentage change in quantity demanded of a commodity due to certain percentage change in its price.


PED = %change in quantity demand/%change in price


Mid point of quantity = "\frac{(Q1 + Q2)} {2}=\frac{ (30 + 20)} {2} = 25"


Mid point of price ="\frac{(P1 + P2)} {2}= \frac{(20 +28)} {2} = 24"


%change in quantity = "\frac{(Q2 - Q1)} {Mid point quantity}= \frac{(20 - 30)} {25} = -0.4"


%change in quantity = "\frac{(P2 - P1)} {Mid point price}= \frac{(28 - 20)} {24} = 0.3"


Price Elasticity of Demand ="\frac{ -0.4} {0.3} = 1.3"


Total Revenue (TR) = Total Quantity "\u00d7" Price per unitTotalQuantity"\u00d7" Priceperunit


Initial TR = "Q_1\u00d7P_1 = 30 \u00d720 = 600"


Present TR = "Q_2\u00d7P_2 = 20 \u00d7 28 = 560"


The total revenue of the firm dropped from $600 to $560